An electron is moving in a circle of radius 12 cm around a magnetic field of T. What is the speed of the electron?
Answer #1

speed of electron in magnetic field is given by cyclotron formula as

v=\frac{qBr}{m}

So, mass m and charge q of electron is known to us ,and r = 12cm = 0.12 m

and B value is missing in question So i take C

so,

v=\frac{1.6\times 10^{-19}\times C\times 0.12}{9.1\times 10^{-31}}

so we get

v=2.1 \times 10^{10}C

So plug in value of magnetic field in place of C to get speed of electron